### v1

@@ 3 Revision theory in standard quantum logic
  • 1.
    $F(\top,\vdash)=\vdash$;
  • 2.
    $F(\alpha,\vdash)=0\quad iff\quad\vdash\neg\alpha$;
  • 3.
    $F(\beta,F(\alpha,\vdash))=F(\alpha,\vdash)\quad iff\quad\alpha\vdash\beta$.
@@
This approach is to consider the operations of property lattice as purely mathematical, which ignores the physical meaning of the operations. For equation $a\wedge_{s}b=a\wedge(\neg a\vee b)=c$, this approach treats the variables $a,b,c$ as indistinctive variables. However, when considering the physical interpretation of operation $\wedge_{s}$, the equation $a\wedge_{s}b=c$ should be interpreted as taking space $a$ as a projection space, projecting space $b$ onto space $a$, and obtaining space $c$ as the result. Moreover, Bob Coecke showed that the Sasaki hook is not a static implicative connective but induces a dynamic one with a parameter. Define $\varphi_{a}^{*}(b):=a\wedge_{s}b$ and $\varphi_{a,*}(b):=a\rightsquigarrow_{s}b$, and then $\varphi_{a}^{*}(b)\leq c\Leftrightarrow b\leq\varphi_{a,*}(c)\Leftrightarrow b\stackrel{{\scriptstyle\varphi_{a}}}{{\rightsquigarrow}}c$ is obtained, where $b$ and $c$ are variables but $a$ is a parameter, which means they are elements from different levels. As external operations of property lattice $L$, the projection operation and its adjoint do not carry out inside a property lattice, but between two property lattices $L_{1}$ and $L_{2}$. Specifically, the element $b$ assigned true in property lattice $L_{1}$ is acted upon by a projection operation $\varphi_{a}$ corresponding to space $a$, resulting in the element $c$ in property lattice $L_{2}$ being assigned true. Therefore, the consequence relation system(CRS) with internalizing connective based on standard quantum logic may not reflect the features of truth-value operations in quantum logic.
@@
4 Truth assignment in topos quantum theory
According to the K-S theorem, all physical quantities cannot be assigned simultaneously and satisfy the functional relation between physical quantities. Isham et al. indicate that contextuality is the core concept to depict a quantum system. Investigating the assignment of physical quantities should examine the assignment of quantities in all contexts. Hence Isham et al. intend to characterize quantum systems with topos theory and inspect the truth-value assignment therein.
4.1 Bi-Heyting algebra structure
@@
Doering shows that the mapping $\underline{\delta}^{o}$ preserves all joins and is an order-preserving injection, but not a surjection. It is evident that $\underline{\delta^{o}(\hat{0})}=\underline{0}$, the empty subobject, and $\underline{\delta^{o}(\hat{1})}=\underline{\Sigma}$. For joins, we have
@@
Doering argues that the elements of $Sub_{cl}(\underline{\Sigma})$ form a complete bi-Heyting algebra. Let $\underline{S},\underline{T}\in Sub_{cl}(\underline{\Sigma})$ be two clopen subobjects. We have the following equations:
@@
$\underline{S}\wedge\underline{R}\leq\underline{T}\quad iff\quad\underline{R}\leq(\underline{S}\rightarrow\underline{T}).$
@@
$(\underline{S}\rightarrow\underline{T})_{V}=\{\lambda\in\underline{\Sigma}_{V}|\forall V^{\prime}\subseteq V\colon if\ \lambda|_{V^{\prime}}\in\underline{S}_{V^{\prime}},then\ \lambda|_{V^{\prime}}\in\underline{T}_{V^{\prime}}\}.$
@@
We have slightly modified Doering’s definition of the operations on co-Heyting algebra. For any $\underline{S}\in Sub_{cl}(\underline{\Sigma})$, the functor
@@
$\underline{T}\leftarrow\underline{S}\leq\underline{R}\quad iff\quad\underline{T}\leq(\underline{R}\vee\underline{S}),$
@@
Theorem 1.
The definition of Heyting implication connective is too trivial for truth-value reasoning, while the definition of co-Heyting implication connective is suitable.
@@
We first examine the definition of Heyting negation $\neg\underline{S}\coloneq(\underline{S}\rightarrow\underline{0})$. Given that $\underline{S}_{V}=\{\lambda_{i}\}_{i\in I}$, we consider $\neg\underline{S}_{V}$. For every $\lambda_{j}\in\underline{\Sigma}_{V}$, if $\lambda_{j}\in\{\lambda_{i}\}_{i\in I}$, then $\lambda_{j}$ does not satisfy the constraint of $\neg\underline{S}_{V}$ in $V$. If $\lambda_{j}\notin\{\lambda_{i}\}_{i\in I}$, then one can always choose a $V^{\prime}\subseteq V$ satisfying $\lambda_{i}|_{V^{\prime}}=\lambda_{j}|_{V^{\prime}}$ such that $\lambda_{j}$ does not satisfy the constraint of $\neg\underline{S}_{V}$ in $V^{\prime}$. Hence $\neg\underline{S}_{V}$ could only be the empty set for every $V$.
@@
Furthermore, we investigate the definition of $\underline{S}\rightarrow\underline{T}$. Without loss of generality, let $\underline{S}_{V}=\{\lambda_{i}\}_{i\in I}$ and $\underline{T}_{V}=\{\lambda_{j}\}_{j\in J}$ (where $I,J$ for index set). For all $\lambda\in\underline{T}_{V}$, $\lambda$ satisfies the constraint of $(\underline{S}\rightarrow\underline{T})_{V}$ obviously. For all $\lambda\notin\underline{T}_{V}$, if $\lambda\in\underline{S}_{V}$, $\lambda$ does not satisfy the constraint of $(\underline{S}\rightarrow\underline{T})_{V}$ in $V$. If $\lambda\notin\underline{S}_{V}$, one can always find a $V^{\prime}\subseteq V$ and some $\lambda_{i}\in\underline{S}_{V}$ satisfying $\lambda|_{V^{\prime}}=\lambda_{i}|_{V^{\prime}}$ such that $\lambda$ does not satisfy the constraint of $(\underline{S}\rightarrow\underline{T})_{V}$ in $V^{\prime}$. Therefore $(\underline{S}\rightarrow\underline{T})_{V}=\underline{T}_{V}$ for every $V$.
Meanwhile, consider the definition of co-Heyting negation
@@
Similarly, consider the definition of co-Heyting implication $(\underline{T}\leftarrow\underline{S})$. Suppose $\underline{S}_{V}=\{\lambda_{i}\}_{i\in I}$ and $\underline{T}_{V}=\{\lambda_{j}\}_{j\in J}$. Take $\{\lambda_{k}\}_{k\in K}=\{\lambda_{j}\}_{j\in J}\backslash\{\lambda_{i}\}_{i\in I}$, and we have $\{\lambda_{j}\}|_{V^{\prime}}\subseteq\{\lambda_{k}\}|_{V^{\prime}}\cup\{\lambda_{i}\}|_{V^{\prime}}$ for every $V^{\prime}\subseteq V$. Hence $(\underline{T}\leftarrow\underline{S})=\bigwedge\{\underline{R}\in Sub_{cl}(\underline{\Sigma})|\forall V,\ \underline{T}_{V}\backslash\underline{S}_{V}\subseteq\underline{R}_{V}\}$. ∎
@@
Given that $\underline{S}_{V}\wedge\underline{R}_{V}\subseteq\underline{T}_{V}$ for some $V$. When considering some $V^{\prime}\subseteq V$, $\underline{R}_{V^{\prime}}$ restricts some elements in $\underline{S}_{V}$ but not in $\underline{T}_{V}$ into $\underline{R}_{V^{\prime}}$. Thus we may have $\underline{S}_{V^{\prime}}\wedge\underline{R}_{V^{\prime}}\not\subseteq\underline{T}_{V^{\prime}}$ in some $V^{\prime}$. Therefore, the set $\underline{R}_{V}$ has to be taken “smaller” and be small enough to satisfy the partial order after being amplified by any $V^{\prime}$. It must be $\underline{R}=\underline{T}$ eventually.
Correspondingly, given that $\underline{T}_{V}\subseteq\underline{R}_{V}\vee\underline{S}_{V}$ for some $V$, it is easy to verify that the partial order is satisfied for any $V^{\prime}\subseteq V$. This is because the elements in $Sub_{cl}(\underline{\Sigma})$ preserve order and join.
Consider the Heyting algebra constituted by elements in $Sub_{cl}(\underline{\Sigma})$. The implicative connective $(\underline{S}\rightarrow\underline{T})=\underline{T}$ is too trivial to apply to truth-value reasoning in the sense of Heyting algebra. But truth-value reasoning in the sense of co-Heyting algebra is suitable. We notice that in every context, the outer daseinisations of elements of orthogonal complement $\hat{P},\hat{P}^{\bot}$ in property lattice $L$ satisfy the law of excluded middle but do not satisfy the law of non-contradiction.
@@
4.2 Heyting algebra structure
In our perspective, for any projection $\hat{P}$, the role of daseinisation is to find the “approximation” of $\hat{P}$ in every context. For a given context $V$, if $\hat{P}\in V$, the approximation of $\hat{P}$ in the context is $\hat{P}$ itself. If $\hat{P}\notin V$, then the approximation of $\hat{P}$ is the element closest to $\hat{P}$ in the context. The outer daseinisation is the element just larger than $\hat{P}$, namely the one that can just be implied by $\hat{P}$. Similarly, the inner daseinisation is the element just smaller than $\hat{P}$, namely the one that can just imply $\hat{P}$. We notice that elements selected by the outer daseinisation in every context satisfy the constraints of morphisms in topos $Set^{\mathcal{V}(\mathcal{H})^{op}}$. Therefore, we regard the outer daseinisation of $\hat{P}$ as a “reflection” of $\hat{P}$ in every context.
Proposition 1.
@@ 4.2 Heyting algebra structure
Theorem 2.
The elements of $D(L)$ constitute a complete Heyting algebra, referred to as the propositional lattice denoted by $H$. Each $\overline{S}\in D(L)$ is called a “truth-value global element”.
@@
5 Natural revision in topos quantum logic
We propose that truth-value reasoning in quantum logic occurs both within a single propositional lattice, as well as between different propositional lattices. For internal truth-value reasoning of an individual element of a propositional lattice, we use the adjoint operation of Heyting algebra. This type of truth-value reasoning that occurs within one propositional lattice is referred to as static quantum logic. The standard quantum logic mentioned in this paper refers to the truth-value reasoning that occurs between two propositional lattices, using external information (projection measurement action) as a parameter to map elements assigned to true of one lattice to those of another. That is, the true-valued elements in the first lattice are mapped to those in the second lattice through Sasaki adjoint operations that use projection space as a parameter. This type of truth-value reasoning that occurs between two propositional lattices is referred to as dynamic quantum logic.
A consequence relation $\vdash\subseteq Fml\times Fml$ is determined by a binary relation obtained from antecedents and consequents. In a propositional lattice, we take the elements that are assigned to true as the antecedents of a consequence relation, and the elements that are inferred to be true according to the partial order relation based on the true-valued elements as the consequents. A truth-value assignment in the propositional lattice corresponds to the antecedents that are associated with a consequence relation. By adding new information, the antecedents of a consequence relation are changed, which is equivalent to revising one consequence relation to another.
When examining natural revision based on quantum logic, we should consider operations in the two types of reasoning - dynamic and static - as two revision operators for consequence relations. We suppose that a consequence relation $\vdash\subseteq Fml\times Fml$ occurs within a propositional lattice, where the formula set $Fml$ represents elements in the propositional lattice, i.e., $Fml=H$. There are two revision operators for revising the consequence relation: one is to add new information using the truth-value reasoning of static quantum logic, and the other is to add new information using that of dynamic quantum logic.
Lemma 1.
The $\wedge$ operation within a Heyting algebra can be seen as a type of revision operation, referred to as object-level revision. Every formula $\alpha\in H$ induces an operator on $\mathcal{C}$

### v2

@@ 3. Revision theory in standard quantum logic
  • $F(\top,\vdash)=\vdash$
  • $F(\alpha,\vdash)=0\quad\text{iff}\quad\vdash\neg\alpha$
  • $F(\beta,F(\alpha,\vdash))=F(\alpha,\vdash)\quad\text{iff}\quad\alpha\vdash\beta$.
The term “consistent consequence relation” refers to a consequence relation that does not yield the contradiction represented by the formula “0”.
@@
This approach is to consider the operations of property lattice as purely mathematical, which ignores the physical meaning of the operations. For equation $a\wedge_{s}b=a\wedge(\neg a\vee b)=c$, this approach treats the variables $a,b,c$ as indistinctive variables. However, when considering the physical interpretation of operation $\wedge_{s}$, the equation $a\wedge_{s}b=c$ should be interpreted as taking space $a$ as a projection space, projecting space $b$ onto space $a$, and obtaining space $c$ as the result. Moreover, Bob Coecke showed that the Sasaki hook is not a static implication connective but induces a dynamic one with a parameter. Define $\varphi_{a}^{*}(b):=a\wedge_{s}b$ and $\varphi_{a,*}(b):=a\rightsquigarrow_{s}b$, and then $\varphi_{a}^{*}(b)\leq c\Leftrightarrow b\leq\varphi_{a,*}(c)\Leftrightarrow b\stackrel{{\scriptstyle\varphi_{a}}}{{\rightsquigarrow}}c$ is obtained, where $b$ and $c$ are variables but $a$ is a parameter, which means they are elements from different levels. As external operations of property lattice $L$, the projection operation and its adjoint do not carry out inside a property lattice, but between two property lattices $L_{1}$ and $L_{2}$. Specifically, the element $b$ assigned true in property lattice $L_{1}$ is acted upon by a projection operation $\varphi_{a}$ corresponding to space $a$, resulting in the element $c$ in property lattice $L_{2}$ being assigned true. Therefore, the consequence relation system(CRS) with internalizing connective based on standard quantum logic may not reflect the features of truth-value operations in quantum logic.
@@
4. Truth assignment in topos quantum theory
According to the Kochen-Specker theorem[20], all physical quantities cannot be assigned simultaneously and satisfy the functional relation between physical quantities. Isham et al. indicate that contextuality is the core concept to depict a quantum system. Investigating the assignment of physical quantities should examine the assignment of quantities in all contexts. Hence Isham et al. intend to characterize quantum systems with topos theory and inspect the truth-value assignment therein.
4.1. Bi-Heyting algebra structure
@@
Döring shows that the mapping $\underline{\delta}^{o}$ preserves all joins and is an order-preserving injection, but not a surjection. It is evident that $\underline{\delta^{o}(\hat{0})}=\underline{0}$, the empty subobject, and $\underline{\delta^{o}(\hat{1})}=\underline{\Sigma}$. For joins, we have
@@
Döring argues that the elements of $Sub_{cl}(\underline{\Sigma})$ form a complete bi-Heyting algebra. Let $\underline{S},\underline{T}\in Sub_{cl}(\underline{\Sigma})$ be two clopen subobjects. We have the following equations:
@@
$\underline{S}\wedge\underline{R}\leq\underline{T}\quad\text{iff}\quad\underline{R}\leq(\underline{S}\rightarrow\underline{T}).$
@@
$(\underline{S}\rightarrow\underline{T})_{V}=\{\lambda\in\underline{\Sigma}_{V}|\forall V^{\prime}\subseteq V\colon\text{if}\ \lambda|_{V^{\prime}}\in\underline{S}_{V^{\prime}},\text{then}\ \lambda|_{V^{\prime}}\in\underline{T}_{V^{\prime}}\}.$
@@
To better align with the adjunction formalism, we have slightly modified Döring’s definition of the operations on co-Heyting algebra. For any $\underline{S}\in Sub_{cl}(\underline{\Sigma})$, the functor
@@
$\underline{T}\leftarrow\underline{S}\leq\underline{R}\quad\text{iff}\quad\underline{T}\leq(\underline{R}\vee\underline{S}),$
@@
Theorem 4.2.
The definition of the Heyting implication connective is considered trivial for truth-value reasoning, whereas the definition of the co-Heyting implication connective presents non-trivial aspects.
Here, triviality refers to the failure of the implication connective to yield effective information that would facilitate logical reasoning.
@@
We first examine the definition of Heyting implication connective $\underline{S}\rightarrow\underline{T}$. The proof is conducted under the following cases.
Case 1: The value of $\underline{S}$ is the empty set in a particular context $V$, denoted as $\underline{S}_{V}=\emptyset$. In this case, $\underline{S}_{V}\rightarrow\underline{T}_{V}=\underline{0}_{V}\rightarrow\underline{T}_{V}=\underline{1}_{V}$. Specifically, if $\forall V,\underline{S}_{V}=\emptyset$ or equivalently $\underline{S}=\underline{0}$, then $\underline{S}\rightarrow\underline{T}=\underline{0}\rightarrow\underline{T}=\underline{1}$.
Case 2: The value of $\underline{T}$ is the empty set in a particular context $V$ but that of $\underline{S}$ is not, denoted as $\underline{T}_{V}=\emptyset$. In this case, $\underline{S}_{V}\rightarrow\underline{T}_{V}=\underline{S}_{V}\rightarrow\underline{0}_{V}=\underline{0}_{V}=\underline{T}_{V}$. Specifically, the value of $\underline{T}$ is the empty set in all contexts, while that of $\underline{S}$ is not always the empty set. That is, $\forall V,\underline{T}_{V}=\emptyset$ or equivalently $\underline{T}=\underline{0}$. Under this setting, $\underline{S}\rightarrow\underline{T}=\underline{S}\rightarrow\underline{0}=\underline{0}=\underline{T}$.
Case 3: The value of $\underline{S}$ in all contexts is a subset of $\underline{T}$, expressed as $\forall V,\underline{S}_{V}\subseteq\underline{T}_{V}$. Consequently, $\underline{S}\leq\underline{T}$, and it is evident that $\underline{S}\rightarrow\underline{T}=\underline{1}$ holds.
Case 4: Consider the values of $\underline{S}$ and $\underline{T}$ in the minimal context $V_{min}$, where the minimal context has the form of $V_{min}=\{0,\hat{P},\hat{P}^{\bot},1\}$. Excluding the four cases where $\underline{S}$ and $\underline{T}$ respectively take $\underline{0}$(or $\underline{1}$) and $\underline{1}$(or $\underline{0}$), which are included in the cases mentioned above, assume $\underline{S}_{V_{min}}=\{\lambda_{\hat{P}}\}$ and $\underline{T}_{V_{min}}=\{\lambda_{\hat{P}^{\bot}}\}$. In this case, it is evident that $(\underline{S}\rightarrow\underline{T})_{V_{min}}=\{\lambda_{\hat{P}^{\bot}}\}=\underline{T}_{V_{min}}$.
@@
Case 5: The general case. Consider the values of $\underline{S}$ and $\underline{T}$ in a non-minimal context $V$, where $\underline{S}_{V}$ is not a subset of $\underline{T}_{V}$. Without loss of generality, let $\underline{S}_{V}=\{\lambda_{i}\}_{i\in I}$ and $\underline{T}_{V}=\{\lambda_{j}\}_{j\in J}$ (where $I,J$ for index set). For all $\lambda\in\underline{T}_{V}$, $\lambda$ satisfies the constraint of $(\underline{S}\rightarrow\underline{T})_{V}$ obviously. For all $\lambda\notin\underline{T}_{V}$, if $\lambda\in\underline{S}_{V}$, $\lambda$ does not satisfy the constraint of $(\underline{S}\rightarrow\underline{T})_{V}$ in $V$. If $\lambda\notin\underline{S}_{V}$, one can always find a $V^{\prime}\subseteq V$ and some $\lambda_{i}\in\underline{S}_{V}$ satisfying $\lambda|_{V^{\prime}}=\lambda_{i}|_{V^{\prime}}$ such that $\lambda$ does not satisfy the constraint of $(\underline{S}\rightarrow\underline{T})_{V}$ in $V^{\prime}$. Therefore $(\underline{S}\rightarrow\underline{T})_{V}=\underline{T}_{V}$ in $V$.
By the attributes of presheaf structure, it can be understood that if $(\underline{S}\rightarrow\underline{T})_{V}=\underline{1}_{V}$ holds within context $V$, then $(\underline{S}\rightarrow\underline{T})_{V^{\prime}}=\underline{1}_{V^{\prime}}$ holds for all $V^{\prime}\subseteq V$.
In summary, aside from the contexts where $\underline{S}_{V}\subseteq\underline{T}_{V}$ (including $\underline{S}_{V}=\emptyset$), which satisfy $\underline{S}_{V}\rightarrow\underline{T}_{V}=\underline{1}_{V}$, in all other instances, $\underline{S}_{V}\rightarrow\underline{T}_{V}=\underline{T}_{V}$. In other words, within any given context $V$, the outcome of $(\underline{S}\rightarrow\underline{T})_{V}$ can only be either $\underline{T}_{V}$ or $\underline{1}_{V}$. Moreover, once $(\underline{S}\rightarrow\underline{T})_{V}=\underline{1}_{V}$ holds in a certain context $V_{1}$, then for all $V_{1}^{{}^{\prime}}\subseteq V_{1}$, it holds that $\underline{S}_{V_{1}^{{}^{\prime}}}\rightarrow\underline{T}_{V_{1}^{{}^{\prime}}}=\underline{1}_{V_{1}^{{}^{\prime}}}$.
Meanwhile, consider the definition of co-Heyting implication $(\underline{T}\leftarrow\underline{S})$. The general case will suffice. Generally, suppose $\underline{S}_{V}=\{\lambda_{i}\}_{i\in I}$ and $\underline{T}_{V}=\{\lambda_{j}\}_{j\in J}$. Take $\{\lambda_{k}\}_{k\in K}=\{\lambda_{j}\}_{j\in J}\backslash\{\lambda_{i}\}_{i\in I}$, and we have $\{\lambda_{j}\}|_{V^{\prime}}\subseteq\{\lambda_{k}\}|_{V^{\prime}}\cup\{\lambda_{i}\}|_{V^{\prime}}$ for every $V^{\prime}\subseteq V$. Hence $(\underline{T}\leftarrow\underline{S})=\bigwedge\{\underline{R}\in Sub_{cl}(\underline{\Sigma})|\forall V,\ \underline{T}_{V}\backslash\underline{S}_{V}\subseteq\underline{R}_{V}\}$. ∎
Specifically, we can examine the definition of Heyting negation $\neg\underline{S}\coloneq(\underline{S}\rightarrow\underline{0})$. Given that $\underline{S}_{V}=\{\lambda_{i}\}_{i\in I}$ for some $V$, we consider $\neg\underline{S}_{V}$. For every $\lambda_{j}\in\underline{\Sigma}_{V}$, if $\lambda_{j}\in\{\lambda_{i}\}_{i\in I}$, then $\lambda_{j}$ does not satisfy the constraint of $\neg\underline{S}_{V}$ in $V$. If $\lambda_{j}\notin\{\lambda_{i}\}_{i\in I}$, then one can always choose a $V^{\prime}\subseteq V$ satisfying $\lambda_{i}|_{V^{\prime}}=\lambda_{j}|_{V^{\prime}}$ such that $\lambda_{j}$ does not satisfy the constraint of $\neg\underline{S}_{V}$ in $V^{\prime}$. Hence $\neg\underline{S}_{V}$ could only be the empty set in $V$.
In the above discussion, two premises are involved: firstly, $\underline{S}$ must select at least one $\lambda$ within the context $V$, and secondly, $V$ must possess a non-trivial subcontext. Addressing these two premises, we examine two special cases. For the first, consider $\underline{S}=\underline{0}$, where $\underline{0}$ does not select any $\lambda$ within $V$, thereby precluding further discussion. It is evident that $\neg\underline{0}=\underline{1}$. For the second case, it is well-known that when considering all subalgebras (i.e., contexts) of $\mathcal{V}(\mathcal{H})$, the context $V=\{0,1\}$ is excluded. Thus, the minimal context in our discussion is in the form of $V_{min}=\{0,\hat{P},\hat{P}^{\bot},1\}$. Consider $\underline{S}$ selecting $\lambda$ in $V_{min}$. Suppose $\underline{S}_{V_{min}}=\{\lambda_{\hat{P}}\}$, then $\neg\underline{S}_{V_{min}}=\{\lambda_{\hat{P}^{\bot}}\}$, and no smaller context $V^{\prime}$ than $V_{min}$ could lead to a contradiction, which indicates $\neg\underline{S}\neq\underline{0}$. Therefore, aside from these two special cases, for all other situations, the discussion holds true, namely, $\neg\underline{S}=\underline{0}$.
Similarly, consider the definition of co-Heyting negation
@@
4.2. Heyting algebra structure
In our perspective, for any projection $\hat{P}$, the role of daseinisation is to find the “approximation” of $\hat{P}$ in every context. For a given context $V$, if $\hat{P}\in V$, the approximation of $\hat{P}$ in the context is $\hat{P}$ itself. If $\hat{P}\notin V$, then the approximation of $\hat{P}$ is the element closest to $\hat{P}$ in the context. The outer daseinisation of $\hat{P}$ is the element just larger than $\hat{P}$, while the inner daseinisation is the element just smaller than $\hat{P}$. We regard the daseinisation of $\hat{P}$ as an “image” of $\hat{P}$ in every context.
Proposition 4.3.
@@ 4.2. Heyting algebra structure
It should be noted that we can utilize the conclusion of Proposition 4.4 to prove Proposition 4.3. From Proposition 4.4, we have $\delta^{i}(\hat{P})\leq\hat{P}\leq\delta^{o}(\hat{P})$. By orthogonal complement, it follows that $\hat{P}^{\bot}\leq\delta^{i}(\hat{P})^{\bot}$ holds for every context. Furthermore, since $\delta^{o}(\hat{P}^{\bot})$ represents the least upper bound of $\hat{P}^{\bot}$ across every context, we obtain $\delta^{o}(\hat{P}^{\bot})\leq\delta^{i}(\hat{P})^{\bot}$ for every context. Consequently, it satisfies $\delta^{i}(\hat{P})\wedge\delta^{o}(\hat{P}^{\bot})\leq\delta^{i}(\hat{P})\wedge\delta^{i}(\hat{P})^{\bot}=0$. Therefore, $\delta^{i}(\hat{P})$ is orthogonal to $\delta^{o}(\hat{P}^{\bot})$, which implies that Proposition 4.3 holds.
@@
$\overline{S}\wedge\overline{R}\leq\overline{T}\quad\text{iff}\quad\overline{R}\leq(\overline{S}\to\overline{T}).$
@@
Theorem 4.7.
The elements of $D(L)$ constitute a complete Heyting algebra.
@@
5. Natural revision in topos quantum logic
We propose a reasoning system that combines the characteristics of two reasoning modes, static and dynamic, as follows: static intuitionistic quantum logic is used as the reasoning rules, and dynamic operations induced by actions are used for truth value assignments.
As Coecke et al.[4, 5, 6] have demonstrated, a projection measurement is an external operation, denoted as $\varphi_{a}^{*}$, parameterized by the projector $a$. This operation acts upon elements within the property lattice, leading to new properties. Moreover, this operation can be extended to the Heyting algebra generated by property sets. When the projection measurement operation is applied to elements in the Heyting algebra that are assigned true (corresponding to the downset of some properties), it can alter the truth value assignments of elements within the Heyting algebra, resulting in new elements being assigned true. For the sake of simplification and provided that it does not lead to any ambiguity, we do not differentiate between the symbols of the $\varphi_{a}^{*}$ operation when applied to either the property lattice or the Heyting algebra.
We use the contextuality to map properties to the corresponding elements in the Heyting algebra and perform the appropriate truth value reasoning. Regarding the changes in truth value assignments caused by external actions, elements within the Heyting algebra that were assigned true in the preceding moment are mapped to elements assigned true in the subsequent moment according to the operational rules of projection measurements.
A consequence relation $\vdash\subseteq Fml\times Fml$ is determined by a binary relation obtained from antecedents and consequents. In Heyting algebra $D(L)$, we take the elements that are assigned to true as the antecedents of a consequence relation, and the elements that are inferred to be true according to the partial order relation based on the true-valued elements as the consequents. A truth-value assignment in the Heyting algebra corresponds to the antecedents that are associated with a consequence relation. By adding new messages, the antecedents of a consequence relation are changed, which is equivalent to revising one consequence relation to another.
When examining natural revision based on quantum logic, we should consider operations in the two types of reasoning - static and dynamic - as two revision operators for consequence relations. We suppose that a consequence relation $\vdash\subseteq Fml\times Fml$ occurs within Heyting algebra $D(L)$, where the formula set $Fml$ represents elements in the Heyting algebra, i.e., $Fml=D(L)$. There are two revision operators for revising the consequence relation: one is to add new message using the truth-value reasoning of static quantum logic, and the other is to add new message by dynamic changes of truth-value assignments.
Lemma 5.1.
The $\wedge$ operation within a Heyting algebra can be seen as a type of revision operation, referred to as static revision. Every formula $\alpha\in D(L)$ induces an operator on $\mathcal{C}$